OpenStudy (anonymous):

Help: please: (9^62773 + 2)^83721= 83721 x log (9^62773 + 2)

OpenStudy (anonymous):

@saifoo.khan help:

OpenStudy (saifoo.khan):

\[\Large (9^{62773} + 2)^{83721}= 83721 \times \log (9^{62773} + 2)\]Like this?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I got 2 but I'm not sure if it's right or not

OpenStudy (saifoo.khan):

There's no variable to solve for! D:

OpenStudy (anonymous):

Imma copy and paste what I got

OpenStudy (saifoo.khan):

Sure.

OpenStudy (anonymous):

log (9^62773 + 2) = log (9^62773) +log ( 2) log (9^62773 + 2) = 62773 log (9) + log ( 2) log (9^62773 + 2) = 62773 (0.954242509) + ( 0.301029996) log (9^62773 + 2) = 59900.966 So: 83721 x log (9^62773 + 2)= (83721)(59900.966) 83721 x log (9^62773 + 2) =5.01496877 × 10^9 Digital sum is 5 + 0+ 1+4+9+6+8+7+7= 47 4+7=11 1+1=2

OpenStudy (anonymous):

@saifoo.khan

OpenStudy (saifoo.khan):

Sorry, i don't understand it. :/